OpenStudy (anonymous):

Calculus: A jumbo-size can of baked beans has a volume of 600cm^3. What dimensions will allow for the minimum amount of metal to produce the can?

OpenStudy (anonymous):

this is a pain but a standard question. i am sure you can look it up and the answer is that the height should equal the diameter

OpenStudy (anonymous):

maybe .sam will write it all out

OpenStudy (anonymous):

\[600=\pi r^2 h\] \[h=\frac{600}{\pi r^2}\] \[A=2\pi r^2 +2\pi rh \] \[A(r)=2\pi r^2 +2\pi r \times\frac{600}{\pi r^2}\] \[A(r) =2\pi r^2 +\frac{1200}{r}\] \[A'(r)=4\pi r -\frac{1200}{r^2}\] set equal zero and solve

OpenStudy (anonymous):

@.Sam. look right?

sam (.sam.):

 \[V=\pi r^2 h\] \[600=\pi r^2h\] \[h=\frac{600}{\pi r^2}\] ======================== \[A=2\pi r(r+h)\] \[A=2\pi r(r+\frac{600}{\pi r^2})\] \[\frac{dA}{dr}=4\pi r-\frac{1200}{r^2}=0\] \[r=300/\pi\] ======================= \[h=\frac{600}{\pi r^2}\] \[h=\frac{600}{\pi (300/\pi)^2}=\frac{1}{150}\pi\]

sam (.sam.):

yep

OpenStudy (anonymous):

nice picture !!

sam (.sam.):

lol thanks